# Locate, Assign, Refine: Taming Customized Image Inpainting with Text-Subject Guidance

### v1

@@ 3.3 Refine Mechanism
$\mathbf{O}_{s}=\text{Softmax}(\frac{\mathbf{Q}_{s}\mathbf{K}_{s}^{\top}}{\sqrt{d}})\mathbf{V}_{s},$ (10)
@@
$\mathbf{Q}_{s}=c_{\text{ctx}}\mathbf{W}_{qs},\mathbf{K}_{s}=[c_{\text{ctx}};c_{\text{obj}}]\mathbf{W}_{ks},\mathbf{V}_{s}=[c_{\text{ctx}};c_{\text{obj}}]\mathbf{W}_{vs},$ (11)
@@
Figure 4: Data construction pipeline. The pipeline involves a sequence of automated processes including categorization, detection, segmentation, and captioning. As a result, it creates region-level quadruplets that include a scene image, a scene mask, a subject image, and a text prompt.
3.4 Training and inference
Data Collection. To our best knowledge, there is no available public dataset that contains both regional text prompts and subject images. So we leverage the power of pretrained large models to produce necessary training data. Specifically, as depicted in Fig. 4, starting from a still image, we first employ RAM [56] to get the included tags within the image. We exclude non-entity tags such as “sky”, “nature” and “skin”. Grounding-DINO [30] and SAM [23] are then utilized in turn to get bounding box and segmentation mask corresponding to each tag. Objects too big or too small are also excluded. Subsequently, we crop the object region according to bounding box and send it to LLaVa [29] to produce the regional caption. By means of this process, we are able to extract quadruplets consisting of (scene image, scene mask, subject image, text prompt) from a vast repository of image data.
Objectives. LAR-Gen is trained with the mean square error loss that is a conventional choice among diffusion-family models. Given the concatenated latent code $\tilde{z}$, the reference subject image $x_{\text{obj}}$ and the text prompt $s$, the loss function is formulated as:
$\mathcal{L}=\mathbb{E}_{\tilde{z},t,x_{\text{obj}},s,\epsilon\in\mathcal{N}(0,\mathbf{I})}\lVert\epsilon-\epsilon_{\theta}(\tilde{z},t,x_{\text{obj}},s)\rVert_{2}$ (12)
Training. A two-stage training strategy is adopted. Specifically, during the first stage, the RefineNet component is omitted, permitting focused training of the main U-Net, augmented by the locate and assign mechanism. The hyper-parameter $\beta$ is currently set to 1. Upon transitioning to the secondary stage, $\beta$ is typically set to a constant value smaller than 1 to ensure the textual semantic alignment, while the RefineNet is integrated to supplement the object detail. At this stage, the main U-Net is frozen, with updates being confined solely to the parameters within the cross-attention layers of the RefineNet.
Inference. We adopt classifier-free guidance[16] during inference, a popular technique that has been proven effective in conditional image generation. Specifically, during training, both the image and text prompt are independently assigned a null value $\emptyset$ with a predetermined probability, thereby enabling the model to concurrently train on both conditional and unconditional denoising capabilities. At inference, it shifts the score estimate towards the conditional direction and away from the unconditional direction:
$\tilde{\epsilon}_{\theta}(\tilde{z},t,x_{\text{obj}},s)=\epsilon_{\theta}(\tilde{z},t,\emptyset,\emptyset)+w(\epsilon_{\theta}(\tilde{z},t,x_{\text{obj}},s)-\epsilon_{\theta}(\tilde{z},t,\emptyset,\emptyset)),$ (13)
@@ 3.4 Training and inference
4 Experiments
@@
Figure 5: Comparison with existing alternatives. Unlike other methods that depend solely on textual or visual prompts, LAR-Gen achieves customized image inpainting by leveraging both the subject image and text prompt.
Table 1: Quantitative results of different approaches on test set. Here $\beta$ represents the image prompt control strength. The condition $s=\emptyset$ indicates that the text condition is omitted, and only the fidelity of the subject is considered.
Method CLIP-I CLIP-T
SD-Inpainting 68.68 30.37
Paint-by-Example 79.02 24.70
Anydoor 82.10 25.52
LAR-Gen ($\beta=0.3$) 79.15 29.94
LAR-Gen ($\beta=1.0,s=\emptyset$) 83.39 26.07
4.1 Experimental Setup
@@
Evaluation Dataset. To evaluate the capability of our method on both subject identity consistency and text semantic consistency, we construct a benchmark that contains 2,000 (scene image, scene mask, subject image, text prompt) samples, using 20 scene images, 10 customized objects, and 10 pre-defined text prompts. The subjects are provided by DreamBooth [38] dataset, and we manually pick 5 non-live subjects and 5 live subjects for fair comparison. For the scene images and scene masks, we randomly pick 20 images with corresponding boxes from the COCO-val dataset [27]. Then, for each scene-subject pair, we bind it with 10 pre-defined prompts to form the final test dataset.
@@
Evaluation Metrics. Following Dreambooth, we evaluate the subject identity consistency by calculating the clip [35] score between the inpainted region and the background-free subject image, denoted as CLIP-I [38]. Additionally, the text semantic consistency is also evaluated by calculating the clip score between the inpainted region and text prompt, denoted as CLIP-T [38].
@@
Implementation Details. We choose Stable Diffusion V1.5 [2] as the base architecture to conduct experiments, while other UNet-based architecture is also compatible with our method, e.g., Stable Diffusion XL. Qualitative results for both architectures are included in the appendix. We use the zoom-in strategy introduced in Anydoor [8], which involves cropping a subimage from the scene image around the masked region to serve as the input for the diffusion model. The model is trained at a batch size of 128 for 20k steps with a learning rate of 1e-5 for both two training stages. The hyper-parameter $\beta$ is set to 1 during the first training stage, and adjusted to 0.3 for the second. The classifier-free guidance scale $w$ is set to 7.5 during inference. All experiments are conducted on A100 gpu of 80G memory.
Figure 6: Ablation studies on varying hyper-parameter $\beta$.
Figure 7: Quantitative analysis of varying the value of hyper-parameter $\beta$. The horizontal axis represents the value of the hyper-parameter $\beta$.
4.2 Comparison with Existing Alternatives
@@
We first compare our method with some existing zero-shot inpainting approaches, including SD-Inpainting [1], Paint-by-Example [49], and Anydoor [8]. Note that SD-Inpainting supports only text-based conditions, whereas Paint-by-Example and Anydoor are limited to using a reference image as the condition. Unlike them, our LAR-Gen is able to inpaint image by simultaneously leveraging a subject image and a text prompt for joint guidance. As depicted in Fig. 5, the qualitative outcomes demonstrate that our LAR-Gen not only ensures that the inpainted region meets the semantic demands of the input prompt but also accurately aligns the generated content with that of the reference image.
@@
We present the quantitative results in Tab. 1. We observe that when $\beta$ is set to 0.3, our method achieves a CLIP-T score comparable to that of SD-Inpainting, greatly surpassing Paint-by-Example and Anydoor, thereby demonstrating its superior capability for maintaining text semantic consistency. Furthermore, our LAR-Gen surpasses SD-Inpainting and Paint-by-Example by 10.47% and 0.13% in terms of CLIP-I score under $\beta=0.3$. However, it trails Anydoor by a margin of 2.94% in terms of CLIP-I scores, due to the fact that the generated object is supposed to be deformed to align the text prompt. We also evaluate the performance of LAR-Gen under the setting of $\beta=1.0$ and no text condition. LAR-Gen achieves the highest CLIP-I score among all approaches, surpassing Anydoor by 1.29%. This result demonstrate its superior capability for preserving the subject identity consistency.
4.3 Ablation Studies
Image Control Strength $\beta$. We conduct ablation experiments to clarify the contribution of hyper-parameter $\beta$. To this end, the first stage model is trained as described previously. Subsequently, in the second stage of training, we adjust $\beta$ from 0.1 to 1.0, which yields a series of second-stage models with varying control strengths. We present the visualization results and quantitative results in Fig. 6 and Fig. 7, respectively. With the increase of scale $\beta$, we observe a greater retention of object details, albeit at the cost of reduced fidelity to the textual descriptions. From Fig. 7 we observe that the CLIP-I score progressively increases with the augmentation of $\beta$, while the CLIP-T score correspondingly diminishes as $\beta$ escalates. This observation corroborates the conclusion that $\beta$ plays a pivotal role in balancing the control between textual semantics and subject identity.
Figure 8: Qualitative ablation studies on each proposed components. Starting with the standard SD-Inpainting model, we incrementally incorporate each component to evaluate its impact on performance.
Table 2: Quantitative ablation studies on core components of LAR-Gen. The Locate mechanism here refers to the use of region-specific captions instead of global captions during training. We assess the performance of Assign mechanism and Refine mechanism by setting $\beta=0.3$.
@@ 4.3 Ablation Studies
Figure 9: Multi-turn image editing results of our method. Thanks to the tailored framework, our model is capable of performing inpainting with various types of guidance. This enables applications that require multi-turn, multi-type inpainting.
Figure 10: Qualitative results on virtual try-on. The cases shown are selected from the test set of VITON-HD.
@@
Effectiveness of Each Component. We conduct experiments to verify the effectiveness of core components proposed in this paper. LAR-Gen consists of three key mechanisms: Locate, Assign and Refine. Starting with the vanilla SD-Inpainting model, we incrementally incorporate each mechanism and assess its impact on text and subject identity consistency by calculating the CLIP-I and CLIP-T scores. In this context, the Locate mechanism specifically refers to the use of regional captions that are associated with particular image regions during training, as opposed to relying on global captions. This is because the SD-Inpainting[1] model already includes the operation of concatenating masked scene images. As shown in Tab. 2, this approach yields a 0.65% improvement in the CLIP-T score over the SD-Inpainting model, suggesting that region-specific captions enhance text semantic consistency due to their more accurate descriptions compared to global captions. Upon this, the addition of the Assign mechanism results in a significant improvement in the CLIP-I score (i.e., 68.60% $\rightarrow$ 75.00%) with only a marginal decrease in the CLIP-T score. This suggests that the Assign mechanism can effectively incorporate the condition from the reference image without substantially compromising text semantic consistency. Lastly, the inclusion of the Refine mechanism leads to a 4.15% increase in CLIP-I score, underscoring its exceptional efficacy in enhancing the high-frequency details of the subject.
4.4 Unified Image Inpainting Framework
@@
Although LAR-Gen is tailored for text-subject-guided image inpainting, it could serve as a unified image inpainting framework that supports text-only and image-only guided inpainting as well. On one hand, to achieve text-only guided inpainting, we set $\beta=0$ and remove RefineNet, thereby our model degenerates to standard SD-Inpainting architecture enhanced by regional caption and performs denoising under text-only guidance. On the other hand, Tab. 1 indicates that by setting the text prompt to $\emptyset$, our model achieves the best fidelity compared to other tuning-free subject-only guided inpainting methods, i.e., Paint-by-Example and Anydoor. It demonstrates the effectiveness of our method on subject-only guided inpainting. Benefiting from the unified framework, our LAR-Gen is capable of performing multi-turn, multi-type inpainting, as depicted in Fig. 9, and show immense potential in a variety of applications, such as virtual try-on, object reshaping, instruct-based image editing, and so on. We present several examples in Fig. 10 to illustrate its capabilities on virtual try-on, which randomly selected from the VITON-HD [9] test set. Note that our model has not been carefully optimized with try-on data. It requires only a single clothing image as input and does not need human pose or any other auxiliary information that generally required on traditional try-on methods. We observe that LAR-Gen achieves seamless cloth composition while preserving fine details of the clothing, such as color, texture, and pattern.
5 Discussion and Conclusion
@@
This paper proposes LAR-Gen, a diffusion-based image inpainting framework that could generate high-fidelity image with the joint guidance of text and subject image. It follows a coarse-to-fine pipeline, which first generates the target subject by introducing the image embedding into the decoupled cross-attention mechanism, then refines the subject details with the proposed RefineNet, avoiding the copy-paste issue caused by collage-based strategy. A hyper-parameter $\beta$ is introduced to balance the strength of image and text condition, thereby enabling fine-grained tuning to seek the optimal setting of preserving both subject identity and text semantic consistency.
@@
Nevertheless, our method still suffers from two limitations. Firstly, the reliance on a single subject image as the conditional input may lead to inaccuracies in subject deformation. This is primarily due to the inherent ambiguity associated with generating deformations that have not been encountered previously, rendering the task inherently ill-posed. Secondly, the denoising process is influenced by multiple conditional factors, which may lead to the neglect of certain conditions, particularly when they conflict with others.
Appendix 0.A Training and Evaluation Data
In this section, we showcase some training samples that are generated using our proposed data construction strategy. This strategy can create quadruplet data consisting of a scene image, a scene mask, a subject image, and a text prompt from a vast repository of image data.  Fig. 11 displays examples of these generated quadruplets. The caption generated for the region accurately details the subject depicted in the subject image, demonstrating its effectiveness in terms of regional semantic alignment.
Additionally, we present the scene and subject images used during the evaluation stage, as shown in Fig. 12. The 20 scene images are randomly selected from COCO-val[27] dataset and the 10 subjects are manually selected from DreamBooth[38] dataset, containing 5 non-live subjects and 5 live subjects. The textual descriptions are also presented in Tab. 3. Following DreamBooth, we design different prompts for non-live and live subjects. Each subject is associated with 10 prompts.
Figure 11: Visualization of some training samples. Note that these are produced by the proposed data construction strategy.
Figure 12: Visualization of scene images and subject images for evaluation. The 20 scene images are randomly selected from COCO-val dataset and the 10 subjects are manually selected from DreamBooth dataset, containing 5 non-live subjects and 5 live subjects.
Table 3: Text prompt list for quantitative evaluation. At inference time, the placeholder S* is replaced with the category label of the specific subject.
Text prompts for non-live subjects Text prompts for live subjects
“a S* on top of a white fabric” “a S* running”
“a S* on top of a purple rug” “a S* on top of a purple rug”
“a S* nearby some books” “a S* nearby some books”
“a S* on top of a wooden box” “a S* wearing a bowtie”
“a red S*” “a S* wearing a top hat”
“a green S*” “a S* plays with a ball”
“a S*, and some sunflowers at around” “a S*, and some sunflowers at around”
“a S*, and some autumn leaves at around” “a S*, and some autumn leaves at around”
“a S* nearby a ball” “a S* wearing sunglasses”
“a S* in front of a cube-shaped metal” “a S* wearing a rainbow scarf”
Appendix 0.B Qualitative Results of Different Backbones
To demonstrate the generalizability of our method, we conduct experiments on Stable Diffusion V1.5[2] and Stable Diffusion XL[4], and present the qualitative results in Fig. 13 and Fig. 14, respectively. The observations indicate that our method performs well on both versions of Stable Diffusion, thereby confirming its effectiveness across different model architectures.
Figure 13: Qualitative results of Stable Diffusion V1.5 backbone.
Figure 14: Qualitative results of Stable Diffusion XL backbone.

### v2

@@ 4.3. Refine Mechanism
(9) $\mathbf{O}_{s}=\text{Softmax}(\frac{\mathbf{Q}_{s}\mathbf{K}_{s}^{\top}}{\sqrt{d}})\mathbf{V}_{s},$
@@
(10) $\mathbf{Q}_{s}=c_{\text{ctx}}\mathbf{W}_{qs},\mathbf{K}_{s}=[c_{\text{ctx}};c_{\text{obj}}]\mathbf{W}_{ks},\mathbf{V}_{s}=[c_{\text{ctx}};c_{\text{obj}}]\mathbf{W}_{vs},$
@@
Figure 3. Data engine, involving a sequence of automated processes including tagging, localizing, segmenting, and captioning.
4.4. Objective and Training
LAR-Gen is trained with the mean square error loss that is a conventional choice among diffusion-family models. Given the concatenated latent code $\tilde{z}$, the reference subject image prompt $x_{\text{obj}}$ and the text prompt $s$, the loss function is formulated as:
(11) $\mathcal{L}=\mathbb{E}_{\tilde{z},t,x_{\text{obj}},s,\epsilon\in\mathcal{N}(0,\mathbf{I})}\lVert\epsilon-\epsilon_{\theta}(\tilde{z},t,x_{\text{obj}},s)\rVert_{2}$
A two-stage training strategy is adopted. Specifically, during the first stage, the RefineNet component is omitted, permitting focused training of the main U-Net, augmented by the locate and assign mechanism. The hyper-parameter $\beta$ is currently set to 1. Upon transitioning to the secondary stage, $\beta$ is typically set to a constant value smaller than 1 to ensure the textual semantic alignment, while the RefineNet is integrated to supplement the object detail. At this stage, the main U-Net is frozen, with updates being confined solely to the parameters within the cross-attention layers of the RefineNet.
4.5. Inference.
We adopt classifier-free guidance(Ho and Salimans, 2021) during inference, a popular technique that has been proven effective in conditional image generation. Specifically, during training, both the image and text prompt are independently assigned a null value $\emptyset$ with a predetermined probability, thereby enabling the model to concurrently train on both conditional and unconditional denoising capabilities. At inference, it shifts the score estimate towards the conditional direction and away from the unconditional direction:
(12) $\tilde{\epsilon}_{\theta}(\tilde{z},t,x_{\text{obj}},s)=\epsilon_{\theta}(\tilde{z},t,\emptyset,\emptyset)+w(\epsilon_{\theta}(\tilde{z},t,x_{\text{obj}},s)-\epsilon_{\theta}(\tilde{z},t,\emptyset,\emptyset)),$
@@ 4.4. Objective and Training
5. Promptable Image Inpainting Data
To the best of our knowledge, there is no public dataset available that contains quadruplet data, including source image, mask prompt, image prompt, and text prompt. So we propose a data engine to automatically generate necessary training data via leveraging the power of pre-trained large models. It consists of four processing stages: tagging, localizing, segmenting, and captioning. Specifically, as depicted in Fig. 3, starting from a still image, we first apply RAM (Zhang et al., 2024) to extract the included tags within the image. We exclude non-entity tags such as “sky”, “nature” and “skin”. Grounding-DINO (Liu et al., 2025) and SAM (Kirillov et al., 2023) are then utilized in turn to localize bounding box and segment mask region of each instance corresponding to each tag. Objects too big or too small are also excluded.
The standard inpainting methods train the model with (image, prompt) pair data, where the prompt is a global description of image. However, we find that global caption may ignore the target subject or fail to describe it in detail. This leads to semantic misalignment on our training stage, and sometimes fail to generate subject. To address this, we crop the object region according to bounding box and send it to LLaVa (Liu et al., 2023) to produce the regional caption. Fig. 4 presents several examples to demonstrate the distinctions between global captions and region-specific captions.
By means of this process, we are able to extract quadruplets consisting of (source image, mask prompt, image prompt, text prompt) from a vast repository of image data, as shown in Fig. 5.
Figure 4. Comparison between global caption and regional caption.
Figure 5. Visualization of some training samples.
6. Experiments
@@
Figure 6. Comparison with existing methods. Our LAR-Gen achieves customized image inpainting corresponding to both text and image prompts, while others depend solely on text or image prompts.
6.1. Dataset and Metrics
@@
To evaluate the capability of our method on both subject identity consistency and text semantic consistency, we construct a benchmark that contains 2,000 (source image, mask prompt, image prompt, text prompt) samples, using 20 scene images, 10 customized objects, and 10 pre-defined text prompts. The image prompts are provided by DreamBooth (Ruiz et al., 2023) dataset, and we manually pick 5 non-live subjects and 5 live subjects for fair comparison. For the source images and mask prompts, we randomly pick 20 images with corresponding bounding boxes from the COCO-val dataset (Lin et al., 2014). The source images and image prompts are shown in Fig. 7. Then, for each (source image, mask prompt, image prompt) pair, we bind it with 10 pre-defined text prompts (as shown in Fig. 8) to form the final test dataset.
@@
Following Dreambooth, we evaluate the subject identity consistency by calculating the clip (Radford et al., 2021) score between the inpainted region and the background-free subject image, denoted as CLIP-I (Ruiz et al., 2023). Additionally, the text semantic consistency is also evaluated by calculating the clip score between the inpainted region and text prompt, denoted as CLIP-T (Ruiz et al., 2023).
Figure 7. Visualization of source images and image prompts for evaluation.
Figure 8. Text prompt list for quantitative evaluation.
Figure 9. Qualitative results of Stable Diffusion V1.5 (left) and Stable Diffusion XL (right).
6.2. Implementation Details
@@
We apply Stable Diffusion V1.5 (AI, 2022b) as the base architecture in our experiments, while other UNet-based architecture is also compatible with our LAR-Gen, e.g., Stable Diffusion XL (AI, 2022c). Qualitative results for both architectures are shown in Fig. 9. We use the zoom-in strategy, which involves cropping a subimage from the source image around the masked region as the input for the diffusion model. The model is trained at a batch size of 128 for 20k steps with a learning rate of 1e-5 for both two training stages. The hyper-parameter $\beta$ is set to 1 during the first training stage, and adjusted to 0.3 for the second. The classifier-free guidance scale $w$ is set to 7.5 during inference. All experiments are conducted on A100 gpu of 80G memory.
6.3. Comparison with Existing Methods
@@
We first compare our LAR-Gen with some existing inpainting methods, including SD-Inpainting (AI, 2022a), Paint-by-Exa (Yang et al., 2023), and Anydoor (Chen et al., 2024a). Note that SD-Inpainting supports only text prompt based condition, whereas Paint-by-Exa and Anydoor are limited to using a reference image prompt as the condition. Unlike them, our LAR-Gen is able to inpaint image by simultaneously leveraging the image and text prompts for joint guidance. As depicted in Fig. 6, the qualitative outcomes demonstrate that our LAR-Gen not only ensures that the inpainted region meets the semantic demands of the input prompt but also accurately aligns the generated content with that of the referenced image prompt.
Table 1. Comparision with SOTA methods. Here $\beta$ represents the image prompt control strength, $s=\emptyset$ indicates that the text prompt is omitted.
Methods Prompts CLIP-I CLIP-T
SD-Inpainting  (AI, 2022a) text&mask 68.68 30.37
Paint-by-Exa  (Yang et al., 2023) image&mask 79.02 24.70
Anydoor  (Chen et al., 2024a) image&mask 82.10 25.52
LAR-Gen ($\beta=0.3$) text&image&mask 79.15 29.94
LAR-Gen ($\beta=1.0,s=\emptyset$) text&image&mask 83.39 26.07
@@
We present the quantitative results in Tab. 1. We observe that when $\beta$ is set to 0.3, our method achieves a CLIP-T score comparable to that of SD-Inpainting, greatly surpassing Paint-by-Exa and Anydoor, thereby demonstrating its superior capability for maintaining semantic consistency with text prompt. Furthermore, our LAR-Gen surpasses SD-Inpainting and Paint-by-Exa by 10.47% and 0.13% in terms of CLIP-I score under $\beta=0.3$. However, it trails Anydoor by a margin of 2.94% in terms of CLIP-I scores, due to the fact that the generated object is supposed to be deformed to align the text prompt. We also evaluate the performance of LAR-Gen under the setting of $\beta=1.0$ and no text condition. LAR-Gen achieves the highest CLIP-I score among all approaches, surpassing Anydoor by 1.29%. This result demonstrates its superior capability for preserving the subject identity consistency with image prompt.
To further verify the superiority of our LAR-Gen, we present a qualitative comparison with SD-Inpainting (AI, 2022a)+IP-Adapter (Ye et al., 2023) (shortened to SDIIPA). The hyper-parameter $\lambda$ is set to 0.5 for SDIIPA (larger $\lambda$ causes the neglect of text) and 0.3 for our method. As shown in Fig. 11, for each pair, the image on the left is generated by SDIIPA, while the right one is by our LAR-Gen. It can be observed that our LAR-Gen consistently outperforms SDIIPA in terms of text semantic alignment and subject identity preservation.
6.4. Ablation Studies
Table 2. Ablation studies on core components of LAR-Gen.
@@ 6.4. Ablation Studies
6.4.1. Effectiveness of Each Component
@@
We conduct experiments to verify the effectiveness of core components in our LAR-Gen, i.e., Locate mechanism, Assign mechanism and Refine mechanism. Starting with the vanilla SD-Inpainting model as baseline, we incrementally incorporate each mechanism and assess its impact on subject identity and text semantic consistencies by calculating the CLIP-I and CLIP-T scores. In this context, Locate mechanism specifically refers to the use of regional captions that are associated with particular image regions during training, as opposed to relying on global captions. This is because the SD-Inpainting(AI, 2022a) model already includes the operation of concatenating masked source images. As shown in Tab. 2, it yields a 0.65% improvement in the CLIP-T score over the SD-Inpainting model, suggesting that region-specific captions enhance text semantic consistency due to their more accurate descriptions compared to global captions. Upon this, the addition of Assign mechanism results in a significant improvement in the CLIP-I score (i.e., 68.60% $\rightarrow$ 75.00%) with only a marginal decrease in the CLIP-T score. This suggests that Assign mechanism can effectively incorporate the condition from the referenced image prompt without substantially compromising text semantic consistency. Lastly, the inclusion of the Refine mechanism leads to a 4.15% increase in CLIP-I score, underscoring its exceptional efficacy in enhancing the high-frequency details of the subject.
Figure 10. Qualitative ablation studies on each proposed components. Starting with the standard SD-Inpainting model, we incrementally incorporate each component to evaluate its impact on performance.
Figure 11. Qualitative results of comparison with IP-Adapter+SD-Inpainting (shortened to SDIIPA). For each pair, the image on the left is generated by SDIIPA, while the right one is by our LAR-Gen.
6.4.2. Image Control Strength $\beta$
We conduct ablation experiments to clarify the contribution of hyper-parameter $\beta$. To this end, the first stage model is trained as described previously. Subsequently, in the second stage of training, we adjust $\beta$ from 0.1 to 1.0, which yields a series of second-stage models with varying control strengths. We present quantitative results in Fig. 12, and visualization results in Fig. 13, respectively. From Fig. 12, we observe that the CLIP-I score progressively increases with the augmentation of $\beta$, while the CLIP-T score correspondingly diminishes. This observation corroborates the conclusion that $\beta$ plays a pivotal role in balancing the control between textual semantics and subject identity. With the increase of scale $\beta$, we observe a greater retention of object details, albeit at the cost of reduced fidelity to the text prompts, as shown in Fig. 13.
Figure 12. Quantitative analysis of varying the value of hyper-parameter $\beta$.
6.5. Diverse Applications
@@
Although LAR-Gen is tailored for multimodal promptable image inpainting, it could serve as a unified image inpainting framework that supports text-only and image-only guided inpainting as well. Benefiting from the unified framework, our LAR-Gen could perform multi-turn, multi-type inpainting, as depicted in Fig. 14, and show immense potential in a variety of applications, such as virtual try-on, object reshaping, instruct-based image editing, and so on. We present several examples in Fig. 16 to illustrate its capabilities on virtual try-on, which randomly selected from the VITON-HD (Choi et al., 2021) test set. Note that our model has not been carefully optimized with try-on data. It requires only a single clothing image as input and does not need human pose or any other auxiliary information that generally required on traditional try-on methods. Our LAR-Gen achieves seamless cloth composition while preserving fine details of the clothing, such as color, texture, and pattern.
Besides, Our LAR-Gen is capable of inpainting parts of objects based on both the subject image and the context. We present some qualitative results to demonstrate this capability, as shown in Fig. 15.
Figure 13. Ablation studies on varying hyper-parameter $\beta$.
Figure 14. Multi-turn and multi-type image editing results of our method.
Figure 15. The ability on inpainting object part.
Figure 16. Qualitative results on virtual try-on. The cases shown are selected from the test set of VITON-HD (Choi et al., 2021).
7. Conclusion
@@
This paper proposes LAR-Gen, a diffusion-based multimodal promptable image inpainting framework that could generate high-fidelity image with the joint guidance of text and subject image. It follows a coarse-to-fine pipeline, which first generates the target subject by introducing the image embedding into the decoupled cross-attention mechanism, then refines the subject details with the proposed RefineNet, avoiding the copy-paste issue caused by collage-based strategy. A hyper-parameter $\beta$ is introduced to balance the strength of image and text condition, thereby enabling fine-grained tuning to seek the optimal setting of preserving both subject identity and text semantic consistency.
@@
Nevertheless, our LAR-Gen still suffers from two limitations. Firstly, the reliance on a single subject image prompt as the conditional input may lead to inaccuracies in subject deformation. This is primarily due to the inherent ambiguity associated with generating deformations that have not been encountered previously, rendering the task inherently ill-posed. Secondly, the denoising process is influenced by multiple conditional factors, which may lead to the neglect of certain conditions, particularly when they conflict with others.